The minimum sum of all possible quantum numbers having n=4 is

Grade:11

2 Answers

Tony
5 years ago
When n=4,
Then 
Minimum value of 4 quantum no. Will be :
 
n=4,
l=0(when n is 4, l can have any value from  0 to 3)
Now, 
Next two quantum numbers has both positive and negative values so use any them  as per question!
m can be - 3,-2,-1,0,1,2,3
s can be - 1/2 or 1/2
Susmita
5 years ago
n=4
l=0,1,2,3
ml=-3,-2,-1,0,1,2,3
ms=+(1/2),-(1/2)
The minimum sum can be achieved by n=4,l=0,ml=0,ms=-(1/2)
Then sun=n+l+ml+ms=4-(1/2)=7/2=3.5
Hope this helps.
Also look at the fact that 3.5 can be achieved by a lot more combinations,eg,n=4,l=3,ml=-3,ms=-(1/2).
But the question only asks for the value of the sum,not the individual quantum numbers.